OpenStudy (anonymous):

A large box contains buttons with pictures of animals. There are four different animals. A button is drawn and replaced. In 100 trials, 35 buttons with cats were drawn. What is the probability that the next button will be a different animal?

OpenStudy (anonymous):

so the probability of a cat is 35/100 hence the probability of it being a different animal would be 100-35/ 100 = 65/100 simplified to 13/20 or if you want it in percentage probability of cat is 35/100=35% so different animal would be 100%-35%=65%

OpenStudy (anonymous):

probability of drawing a cat*

OpenStudy (amistre64):

without knowing the total number of buttons, and how many button have which animal ... id say its impossible to determine. Each new trial is independant of the last one, so the probability of picking 'not a cat' is the same as it was for any of the other trials.

OpenStudy (amistre64):

is 100 times a sufficiently large number to conclude the law of large numbers?

OpenStudy (anonymous):

it changes every time, but at that point it would be 65% the next would likely be ?/101

OpenStudy (amistre64):

if we are basing this off of the experimental trials, then yeah, i agree with 65% .... but i do not see that as a good standard to base the experiment from.

OpenStudy (anonymous):

ah well introduction to probability can't be too complicated